OpenStudy (anonymous):

Lisa and Gage are painting a room. If Lisa has painted 1/6 of the room and Gage has painted 1/4 of the room, what fraction of the room remains to be painted?

OpenStudy (anonymous):

Whenever you are working with fractions make sure you put all the fractions in common denominators! Then add the two fractions and see how much of the room has been painted

OpenStudy (anonymous):

2/3

OpenStudy (anonymous):

actually it is 1/6 + 1/4 = 2/12 + 3/12 = 5/12

OpenStudy (anonymous):

That is what I got Walleye

OpenStudy (anonymous):

But ...

OpenStudy (anonymous):

What fraction of the room REMAINS TO BE PAINTED, therefore 12/12 - 5/12= 7/12

OpenStudy (anonymous):

ah yes good point patbat. You're correct!

OpenStudy (anonymous):

my bad... u r right!

OpenStudy (anonymous):

I'm gonna post some more

OpenStudy (anonymous):

ok ill take a look